# Ultrastable nanophotonic microcavities via integrated thermometry

### v1

@@ Abstract
Integrated photonic devices that can be co-packaged with electronics and can operate in real-world environments will enable many important applications, such as optical interconnects, quantum information processing, precision measurements, spectroscopy, and microwave generation. Significant progress has been made over the past two decades on increasing the functional complexity of photonic chips. However, the most critical challenge that remains is the lack of scalable techniques to overcome perturbations arising from environmental thermal noise and thermal crosstalk from co-packaged electronics and other photonic devices sharing the same substrate. Here, we propose and demonstrate a fully-integrated scheme to monitor and stabilize the temperature of a high-$Q$ microresonator in a Si-based chip. We show that when stabilized, the microresonator exhibits remarkable resilience against external thermal noise and can serve as a fully-integrated photonic frequency reference. By simply changing the temperature set-point, the cavity resonance frequency can be repeatably tuned without any hysteresis over several hours. We also show that the frequency of a distributed feedback (DFB) laser can be stabilized to this microresonator and realize a 48× reduction in its frequency drift, resulting in its center wavelength staying within ±0.5 pm of the mean over the duration of 50 hours in the presence of significant ambient fluctuations. This performance is superior to many commercial DFB systems and is highly suitable for use in data-communication systems. Finally, we demonstrate that the technique can be implemented to stabilize a soliton modelocked Kerr comb against significant ambient and crosstalk thermal noise, without the need for photodetection, paving the way for Kerr-comb-based photonic devices that can operate in the desired modelocked state indefinitely.
@@
I Introduction
Optical frequency references are critically important tools for metrology [1, 2, 3], optical atomic clocks [4, 5, 6], precision spectroscopy [7, 8, 9], and telecommunication systems [10, 11, 12]. At wavelengths away from molecular, atomic, and nuclear transitions, stable optical cavities with narrow linewidths and low long-term drift are indispensable for laser frequency stabilization. A laser’s frequency can be stabilized by measuring its frequency detuning from the resonance of a stable narrow-linewidth cavity, and performing electronic or optical feedback correction. Within the bandwidth of the feedback mechanism, the cavity’s frequency stability can be transferred to the laser. To this end, cavities with large optical mode volumes offer exceptional fundamental frequency stability due to their low Brownian thermorefractive noise limit. For example, bulk Fabry-Perot (FP) cavities that utilize single-crystal Silicon and low-thermal-coefficient mirror coatings and housings operating at cryogenic temperatures have been reported to exhibit a fractional frequency stability of $10^{-17}$ over one second [13, 14]. Despite their exceptional stability, cryogenically-cooled bench-top systems are not well-suited for practical applications that require compactness, portability, and robustness. The rapid development of integrated low-loss dielectric photonic platforms over the past two decades offers the prospect of realizing stable reference microcavities in small form factors [15, 16]. However, the development of stable microcavities is challenging due to the enhanced sensitivity to fundamental and ambient temperature fluctuations that occur with small mode volumes, which in turn leads to significant drift in the cavity’s absolute resonance frequencies [17, 18].
@@ I Introduction
High-$Q$ microcavities have proven to be excellent platforms for the generation of frequency combs, which have a number of applications such as spectroscopy [19, 20, 21], chip-scale multi-wavelength sources [22, 23], microwave generation [24, 25, 26], light detection and ranging (LiDAR) [27, 28], and optical interconnects for large-scale machine learning [29, 30]. Although monolithically-integrated microcavities offer superior mechanical stability than that of bulk cavities, a central challenge that remains is the lack of resource-efficient techniques to stabilize against thermal perturbations arising from ambient sources and thermal crosstalk from other thermally-tuned devices sharing the same substrate [31, 32]. For Kerr-comb applications, small temperature fluctuations can readily lead to pump detunings that fall outside of the modelocked regime of operation. Thermal crosstalk is a key bottleneck for neuromorphic computing applications [32] and densely-integrated co-packaged photonic input/output (I/O) modules [30], where multiple photonic devices on the same chip are typically simultaneously operated.
@@
Passive methods to reduce Brownian thermal noise has been demonstrated previously by leveraging the inverse relationship with the mode volume, using a large-mode-volume integrated spiral cavity [15]. Manipulation of heat-flow, such as using undercuts [33] and trenches [34], have been shown to partially reduce the impact of external thermal perturbations, while adding significantly to fabrication cost and complexity. Active feedback using the photoconductive effect in silicon has been previously reported for stabilizing the temperature of microresonators [35, 36, 37]. Such a method, however, is inherently limited by the phase and intensity noise of the probing laser, and requires the use of highly-stable bench-top laser systems. Additionally, because the method relies upon the photoconductive effect to indirectly infer cavity temperature, its relevance is limited to semiconductor materials, and is not applicable to dielectric platforms that are commonly used in Kerr-comb-related applications. Laser frequency stabilization via dual-mode optical thermometry using a high-$Q$ microresonator has also been reported [38, 16]. However, it requires optical probing involving multiple phase modulators and cumbersome electronics, the excitation of two polarization or frequency modes, as well as exceptionally high quality factors for both two excited modes. All of these drawbacks make such a technique ill-suited for practical applications that require compactness and resource-efficient designs. Finally, for comb-based devices, active-feedback stabilization based on monitoring the optical output has been reported [39, 40]. However, the requirement of additional drop-ports, optical filters, photodiodes, and in many cases, free-space or fiber optics, significantly adds to the footprint and fabrication cost.
@@
Here, we demonstrate a highly general approach, based on fully-integrated thermometry, to stabilize the temperature, and thereby the absolute resonance frequency, of a high-$Q$ microcavity in the presence of significant external noise over $24$ hours. A thin-film Platinum resistor is designed to exhibit a strong temperature-dependent electrical resistance arising from the metal’s intrinsic temperature coefficient of resistivity (TCR). Similar metallic structures have previously been extensively used as microheaters [41, 42, 43, 44]. We utilize such a resistor as an integrated resistance thermometer, enabling real-time monitoring of a microcavity’s temperature without the need for any external photodetection or other integrated non-linear electronic elements (e.g., diodes, transistors). The use of sputtered thin-films in our fabrication eliminates the need for processes requiring high-temperature and/or corrosive processing (e.g., electroplating). Moreover, Platinum being chemically resistant to basic chip cleaning reagents as well as atmospheric humidity leads to device longevity while ensuring long-term repeatability of our measurements with negligible variations. By engineering a dual-resistor scheme that ensures proper thermal stability of the microresonator core, we show that the absolute resonance frequency of a high-$Q$ silicon nitride (SiN) microcavity can be stabilized even in highly noisy environments to within tens of MHz of any desired value and can be repeatably tuned by simply adjusting an electrical voltage. The frequency noise of such a microcavity is shown to be impervious to significant environmental drift and thermal crosstalk, and is only limited by flicker noise in the control electronics. The resonance frequency of the stabilized microresonator exhibits bounded fluctuations over extended durations of time, in contrast to the unbounded random-walk behavior demonstrated by free-running microresonators and compact distributed feedback (DFB) lasers. By locking a DFB laser to the stabilized microresonator, we demonstrate that the microresonator can serve as a fully-integrated photonic frequency reference for applications that require long-term stability in a small form factor while operating in noisy environments, such as in integrated photonic transceivers for dense wavelength-division multiplexing (DWDM) applications. The locked laser exhibits a long-term stability of emission center wavelength of <±$<\pm 0.5$ pm over $50$ hours of measurement duration, significantly better than many commercial systems. Finally, we apply integrated thermometry to demonstrate a fully-thermally-stabilized Kerr frequency comb that remains stably modelocked in the presence of strong thermal crosstalk. We envision this work to enable the realization of highly robust and field-deployable linear and nonlinear photonic devices for classical and quantum applications, and integrated Kerr combs that can operate indefinitely.
II Integrated Thermometry
@@
A conceptual illustration of integrated thermometry is provided in Fig. 1(a). A thin-film metallic resistor with a temperature-dependent resistivity, placed directly above a high-$Q$ microcavity, acts as an on-chip resistance thermometer. As a result of the low heat capacity of the thin-film resistor, small changes in heat flow can translate into large and observable changes in its temperature, and consequently in its electrical resistance. This enables monitoring changes in the microcavity temperature due to external noise, which is especially critical in practical scenarios. The current-voltage-resistance (I-V-R) characteristics of the metallic resistor, experimental measurements of which are shown in Fig. 1(b), indicate a deviation from the linear I-V trend, corresponding to a voltage-dependent resistance. This qualitative behavior is identical to a filament lamp, which exhibits lower (higher) resistance at lower (higher) voltages due to lower (higher) filament temperature arising from Joule heating. The near-quadratic dependence of resistance on applied voltage indicates a linear dependence on the power dissipated. Figure 1(c) shows a similar relationship between the thermometer’s resistance and the voltage applied across the second heater, demonstrating that temperature changes arising from external heat sources can be measured by simply monitoring the electrical resistance of the resistance thermometer.
Figure 1: a, Conceptual illustration of integrated thermometry applied to stabilize a high-Q monolithic microresonator against thermal noise arising from ambient heat sources and crosstalk from other thermally-tuned devices on the same chip. A thin-film metallic (platinum) resistor, typically used as a microheater, exhibits a temperature-dependent resistance due to the metal’s natural temperature-dependent resistivity. As a result of the low heat capacity of the thin-film resistor, small fluctuations in the heat flow in its vicinity lead to large changes in its temperature, enabling the real-time monitoring of on-chip temperature by simply measuring its electrical resistance. A second identical resistor is used as a heater to perform active stabilization purely using the resistance thermometer, thereby eliminating the need for optical probing for stabilization. Image for data center sourced from Ref. [45]. b, Current-voltage-resistance (I-V-R) characteristics of the resistance thermometer fabricated in this work. The I-V measurements (pale blue dots) deviate from the linear trend (dark blue line), and the resistance (yellow dots) exhibits a quadratic dependence (purple line) on the thermometer voltage. c, The measured thermometer resistance (lavender dots) also exhibits a quadratic dependence (green line) on the voltage applied across the (second) heater, indicating a linear dependence on power dissipated across the heater. d, Resonance frequency shift (green circles) of a high-Q microcavity of free-spectral range (FSR) 76 GHz, measured using a piezo-tuned tunable laser source, and measured change in thermometer resistance (gold line), for a sinusoidal perturbation applied via the heater. The cavity temperature and resonance frequency exhibit a strong anti-correlation. e, Measured resonance frequency shift plotted versus temperature measured by the thermometer, same data as in d. The linear fit provides an in-situ measurement of the thermo-optic coefficient of the fundamental spatial mode of the cavity waveguide.
The temperature measurement scheme can be utilized together with optical measurements to perform rapid in-situ measurements of the thermo-optic coefficient of the microcavity waveguide’s spatial mode at any desired wavelength. In order to demonstrate this, a small-amplitude 2-Hz sinusoidal voltage waveform is applied to the heater, while monitoring the electrical resistance of the thermometer. The free-spectral-range (FSR) of the SiN microresonator is 76 GHz, its loaded $Q$ is approximately $3\times 10^{6}$, and the resonances in the telecommunications C-band are critically coupled. The width of the waveguide is 2100 nm. A tunable laser source centered at 1560 nm is swept at 200 Hz, using its internal piezo-electric module, across a single microcavity resonance. By synchronously monitoring the transmitted optical output using a photodiode connected to a fast oscilloscope, the resonance frequency shift of the cavity is measured. Simultaneously, the on-chip temperature change is also monitored on the same oscilloscope by sending a constant low-noise DC current through the resistance thermometer and measuring the change in voltage across it. Figure 1(d) shows the measured resonance frequency shift (green circles; left y-axis) and the corresponding change in thermometer resistance (golden solid line; right y-axis), indicating a clear correlation between the two. The measured change in thermometer resistance is converted to a change in temperature using the following equation,
$R(T)=R_{0}(1+A\Delta T)$ (1)
where $R_{0}$ is the resistance at room temperature, $\Delta T$ is the change in temperature, and $A$ is the temperature coefficient of resistance (TCR). The coefficient $A$ was measured independently to be $A=0.00172$. The measurement procedure is detailed in Methods.
@@
Figure 1(e) shows the data presented in Fig. 1(d), now with the resonance frequency shift plotted against the measured change in temperature. The slope of the linear fit is then used to extract the thermo-optic coefficient of the fundamental spatial mode of the waveguide, $dn_{eff}/dT=2.45\times 10^{-5}$ K-1. This value is also approximately equal to the thermo-optic coefficient $dn/dT$ of SiN due to the high spatial confinement offered by the fundamental mode of the 2100-nm waveguide, and is consistent with values published previously that employ different measurement techniques [46, 47, 48].
III Fully-Integrated Photonic Frequency Reference
@@
The frequency stability of the microresonator is investigated, as depicted in Fig. 2(a), by piezo-scanning a probe laser across a cavity resonance at 200 Hz, and recording the resonance frequency position once every ten seconds. Absolute wavelength calibration is performed by beating the tunable laser with a low-drift fixed-wavelength reference laser on a fast photodiode, and simultaneously calibrating the reference laser’s wavelength using a benchtop wavemeter (Bristol Instruments 238A). This yields a calibrated absolute cavity resonance frequency, which is recorded continually once every 10 seconds over the course of 24 hours. The microcavity’s frequency stability is measured under the following three conditions: (i) ‘free-running’, in which the cavity is subject only to ambient fluctuations (such as change in laboratory temperature), (ii) ‘open-loop’, in which an additional perturbation is induced via a voltage summing circuit connected to the heater, in order to emulate real-world use, and (iii) ‘stabilized’, in which the microresonator’s temperature is actively stabilized. The strength of the perturbation added is denoted by the parameter $\xi$, which is defined as follows. For $\xi$=1, the root mean-squared (RMS) fluctuation of the cavity resonance frequency equals its short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity used in this work. The cavity is stabilized by sensing the temperature in real-time by sending a low-noise electrical current through the resistance thermometer and measuring voltage changes across it, and by performing active feedback to the heater using a low-noise proportional-integral-derivative (PID) circuit. A schematic of the current source fabricated for this work and its noise performance are provided in Supplementary Information. The thermal design of the dual-resistor scheme described here is explained further in Methods.
Figure 2: Long-term stabilization of the absolute resonance frequency of a high-Q SiN microresonator using integrated thermometry. a, Simplified experimental schematic, depicting a tunable probe laser that is piezo-scanned across a microcavity resonance. Absolute frequency calibration of the scan is performed by producing a heterodyne note between the tunable laser and a stable reference laser, and simultaneously tracking the reference laser’s wavelength drift using a highly-stable benchtop wavelength meter. The heterodyne beat note as well as the cavity resonance line-shape are measured synchronously with the piezo sweep using an real-time oscilloscope connected to a computer interface for periodic data acquisition every 10 seconds. For $\xi$=1, the pseudorandom perturbation applied via the outer heater yields a root mean-squared (RMS) of the induced cavity resonance frequency fluctuation equals its short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity fabricated in this work. b, Drift in the microcavity’s calibrated resonance frequency over the duration of 24 hours, for free-running ($\xi$=0), open-loop ($\xi$=7.55), and stabilized ($\xi$=7.55) operation. The free-running cavity exhibits a slow but strong drift in its resonance frequency due to ambient temperature drift in the lab, whereas the open-loop cavity additionally exhibits strong “fast” fluctuations due to the perturbation introduced. The stabilized cavity’s resonance frequency remains highly stable in the presence of very strong ambient and crosstalk-induced perturbations. c, Allan Deviation (ADEV) versus averaging time, computed using the data displayed in b. The stabilized cavity exhibits an ADEV that is well below the free-running and open-loop modes of operation. The shape of the ADEV curve of the stabilized cavity matches that of $1/f$ noise, indicating that performance is only limited by control electronics. d, Histogram plots depicting the fraction of time over 24 hours (y-axis) that the microcavity’s absolute resonance frequency drift lies within a given frequency bin with a width of 5 MHz.
@@
Figure 2(b) displays the measured resonance frequency drift of the microresonator for the three cases described above over a period of 24 hours. In the absence of any stabilization, ambient noise alone causes a significant resonance drift, even in temperature-stabilized laboratory environments. This is shown by the slow cyclical pattern in the ‘free-running cavity’ data points, with an standard deviation of resonance frequency variation of $\approx 341$ MHz. Additional perturbation, such as from thermal crosstalk in practical applications, adds to the drift in the form of faster-timescale fluctuations, as evidenced by the ‘open-loop cavity’ data points, exhibiting a standard deviation of resonance frequency variation of $\approx 470$ MHz. The Supplementary Information provides the corresponding real-time temperatures measured by the resistance thermometer for the free-running and open-loop cases, which shows excellent correlation with the resonance frequency shift. Finally, the stabilized cavity exhibits a highly stable resonance frequency with a considerably lower drift even in the presence of significant added perturbation ($\xi=7.55$). The standard deviation of the measured resonance frequency variation of the stabilized microresonator is $\approx 13.7$ MHz, corresponding to a 34×34$\times$34 × reduction compared to the open-loop scenario. The time-domain measurement data is better visualized as histograms, shown in Fig. 2(d). The bin-size along the x-axis is 5 MHz. Remarkably, even in the absence of additional perturbations, i.e., in the free-running case, the cavity resonance frequency histogram does not consist of a single, well-defined narrow peak.

### v2

# Frequency-stable nanophotonic microcavities via integrated thermometry

@@ Abstract
Field-deployable integrated photonic devices co-packaged with electronics will enable important applications such as optical interconnects, quantum information processing, precision measurements, spectroscopy, and microwave generation. Significant progress has been made over the past two decades on increasing the functional complexity of photonic chips. However, a critical challenge that remains is the lack of scalable techniques to overcome thermal perturbations arising from the environment and co-packaged electronics. Here, we demonstrate a fully integrated scheme to monitor and stabilize the temperature of a high-$Q$ microresonator on a Si-based chip, which can serve as a photonic frequency reference. Our approach relies on a thin-film metallic resistor placed directly above the microcavity, acting as an integrated resistance thermometer, enabling unique mapping of the cavity’s absolute resonance wavelength to the thermometer’s electrical resistance. Following a one-time calibration, the microresonator can be accurately and repeatably tuned to any desired absolute resonance wavelength using thermometry alone with a root-mean squared wavelength error of <0.8<0.8 pm over a timespan of days. We frequency-lock a distributed feedback (DFB) laser to the microresonator and demonstrate a 48× reduction in its frequency drift, resulting in its center wavelength staying within ±0.5 pm of the mean over the duration of 50 hours in the presence of significant ambient fluctuations, outperforming many commercial DFB and wavelength-locker-based laser systems. Finally, we stabilize a soliton mode-locked Kerr comb without the need for photodetection, paving the way for Kerr-comb-based photonic devices that can potentially operate in the desired mode-locked state indefinitely.
@@
Optical frequency references are critically important tools for metrology [1, 2, 3], optical atomic clocks [4, 5, 6], precision spectroscopy [7, 8, 9], and telecommunication systems [10, 11, 12]. At wavelengths away from molecular, atomic, and nuclear transitions, optical cavities with narrow linewidths and low long-term drift are indispensable for laser frequency stabilization. Cavities with large optical mode volumes and low Brownian thermorefractive noise (TRN) offer the best frequency stability, such as bulk single-crystal Silicon Fabry-Perot (FP) cavities with a reported fractional frequency instability of $10^{-17}$ over one second [13, 14]. For many applications that require portability and robustness, such as in telecommunication systems, cryogenically-cooled bulk cavities are not well-suited, necessitating the use of compact reference and laser cavities that trade off size, weight, and power consumption (SWaP) for frequency stability that is typically several orders of magnitude worse compared to state-of-the-art bulk systems. In addition to the plethora of commercially available temperature-stabilized distributed feedback (DFB) and distributed Bragg reflector (DBR) lasers, several compact frequency stabilization technologies have been demonstrated in recent years, including integrated optical frequency discriminators [15, 16], vapor-cell-integrated photonic chips [17], compact wavelength lockers [18], and fiber interferometers [19].
The rapid development of low-loss dielectric photonic platforms over the past two decades offers the prospect of realizing stable reference cavities in small form factors [20, 21]. In recent years, microcavities with large quality factors ($Q$s) have been shown to enable short-term frequency instability approaching $10^{-14}$ [22, 23]. However, achieving long-term stability remains challenging due to the enhanced sensitivity to fundamental and ambient temperature fluctuations [24, 25]. High-$Q$ microcavities are also excellent platforms for the generation of frequency combs, which have a number of applications such as spectroscopy [26, 27, 28], chip-scale multi-wavelength sources [29, 30], microwave generation [31, 32, 33], light detection and ranging (LiDAR) [34, 35], and optical interconnects for large-scale machine learning [36, 37].
@@
A central challenge that remains to be addressed in integrated photonics broadly is the lack of resource-efficient techniques to stabilize against thermal perturbations arising from ambient sources, co-packaged electronics, and thermal crosstalk [38, 39]. For Kerr-comb applications, small temperature fluctuations can readily lead to pump detunings that lie outside of the modelocked regime of operation. Thermal crosstalk is a key bottleneck for neuromorphic computing applications [39] and densely integrated co-packaged photonic input/output (I/O) modules that require simultaneous tuning of multiple on-chip devices [37].
@@
Passive methods to mitigate thermal perturbations have been demonstrated previously, such as large mode volumes [20], and manipulating heat flow using undercuts [40] and trenches [41]. However, passive methods only partially reduce the impact of thermal perturbations while adding significantly to fabrication cost and complexity. Active stabilization of silicon microresonators, based on the photoconductive effect, to a reference laser has been shown [42, 43, 44]. However, such a method is inherently limited by the reference laser’s stability and is not applicable to high-$Q$ dielectric platforms that are commonly used for realizing Kerr combs. Laser frequency stabilization via dual-mode optical thermometry using a high-$Q$ microresonator has also been reported [45, 21]. However, it requires multiple phase modulators, cumbersome electronics, and the excitation of modes with proximate resonance frequencies and very high $Q$s. For comb-based devices, while active feedback using the optical output has been reported [46, 47], the requirement of additional drop-ports, optical filters, photodiodes, and free-space/fiber optics significantly adds to footprint and fabrication cost.
@@
Here, we demonstrate a highly general approach, based on fully integrated thermometry, to stabilize the temperature of a high-$Q$ microcavity in the presence of significant external perturbations. A thin-film Platinum resistor is designed to exhibit a strong temperature-dependent electrical resistance and is used as an integrated resistance thermometer to measure a microcavity’s temperature without the need for any photodetection or other integrated non-linear electronic elements (e.g., diodes, transistors). Similar metallic structures have previously been extensively used as microheaters [48, 49, 50, 51]. Platinum’s chemical stability against basic chip cleaning reagents and atmospheric humidity ensures device longevity and long-term repeatability of our measurements. By engineering a dual-resistor scheme that achieves proper thermal stability of the microresonator waveguide core, we show that the absolute resonance frequency of an unshielded and unpackaged high-$Q$ silicon nitride (SiN) microcavity can be stabilized to within tens of MHz of any desired value and can be tuned repeatably by simply adjusting an electrical voltage. The frequency noise of such a microcavity is shown to be impervious to environmental drift and thermal crosstalk and limited only by $1/f$ flicker noise in the control electronics. We demonstrate that a one-time calibration of the resonance wavelength versus thermometer resistance is sufficient to repeatably tune the cavity to any desired resonance wavelength within the calibration range using thermometric control alone for at least a few days, removing the requirement for optical referencing or monitoring.
By locking a DFB laser to the stabilized microresonator, we demonstrate that it can serve as a fully-integrated photonic frequency reference for telecommunications transceiver applications. The locked laser’s emission wavelength remains within $\pm 0.5$ pm of the mean over $50$ hours of measurement duration, outperforming many commercial DFB and wavelength-locker-based laser systems. Finally, we apply integrated thermometry to demonstrate a fully-thermally-stabilized Kerr comb that remains stably modelocked in the presence of strong thermal crosstalk. We envision this work to enable the realization of highly robust and field-deployable linear and nonlinear photonic devices for classical and quantum applications, and integrated Kerr combs that can potentially operate indefinitely.
I Results
@@
Figure 1(a) conceptually illustrates integrated thermometry. A thin-film metallic resistor with a temperature-dependent resistance placed directly above a high-$Q$ microcavity acts as an on-chip resistance thermometer. Due to the low heat capacity of the thin-film resistor, small heat fluxes translate to large and observable changes in its temperature. The current-voltage-resistance (I-V-R) characteristics displayed in Fig. 1(b) show a non-linear I-V relationship, and a quadratic dependence of R on V, indicating a linear dependence of R on the dissipated power. This behavior is similar to that of a filament lamp, which exhibits lower (higher) resistance at lower (higher) voltages due to lower (higher) filament temperature from Joule heating. Figure 1(c) shows that the thermometer resistance also depends quadratically on the voltage applied across the second heater, demonstrating that temperature changes induced by external heat sources can be measured by simply monitoring the resistance. Extended Data Fig. 2 shows an independent measurement of the temperature-dependent resistance [52].
@@
We combine the temperature measurement scheme described above with a scanning-laser optical probe to perform rapid in-situ measurements of the thermo-optic coefficient of a silicon-nitride (SiN) microresonator’s waveguide mode. The free-spectral range (FSR) is 76 GHz, and the loaded $Q$ is approximately $3\times 10^{6}$. The measurement procedure is described in Methods. Figure 1(d) shows the measured resonance frequency shift and the corresponding change in thermometer resistance, illustrating a strong correlation. Figure 1(e) plots the resonance frequency shift versus the measured temperature change. The slope of the linear fit is used to extract the thermo-optic coefficient of the fundamental spatial mode of the SiN waveguide, $dn_{\mathrm{eff}}/dT=2.45\times 10^{-5}$ K-1. This value is approximately equal to the thermo-optic coefficient $dn/dT$ of SiN due to the high spatial confinement offered by the 2100-nm waveguide and is consistent with values measured using other techniques [53, 54, 55].
@@
Figure 2(a) illustrates the experimental schematic for investigating the long-term frequency stability of the microresonator using a piezo-controlled scanning probe laser. The absolute frequency calibration procedure is described in Methods. The microcavity resonance frequency is averaged over and recorded once every 10 seconds over the period of 24 hours. Three such measurements are performed, each under the following conditions, (i) ‘free-running’: the cavity is subject only to ambient temperature fluctuations, (ii) ‘open-loop’: additional perturbations are induced via a voltage-summing circuit connected to the heater to mimic thermal crosstalk from co-packaged electronics, and (iii) ‘stabilized’: the microresonator’s temperature is actively stabilized by applying proportional-integral-derivative (PID) feedback to the heater based on the voltage drop across the resistance thermometer connected to a low-noise current source. The strength of the perturbation applied is denoted by the empirical parameter $\xi$: for $\xi$=1, the root mean-squared (RMS) fluctuation of the resonance frequency equals the short-term resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity used in this work. A circuit schematic of the current source and its noise performance are provided in Supplementary Fig. S1. The thermal design of the dual-resistor scheme is explained in Extended Data Fig. 1.
@@
Figure 2(b) displays the measured resonance frequency fluctuations. Without stabilization, ambient noise causes a significant resonance drift, even in a temperature-stabilized laboratory, as seen in the ‘free-running cavity’ trace. The resulting standard deviation of the resonance frequency is $\approx 341$ MHz. The additional perturbation applied ($\xi=7.55$) drives faster-timescale fluctuations seen in the ‘open-loop cavity’ trace. The measured open-loop standard deviation is $\approx 470$ MHz. Supplementary Figs. S2 and S3 provide the temperatures measured by the resistance thermometer for the free-running and open-loop cases, which show strong correlations with the resonance frequency shifts. Finally, the stabilized cavity exhibits a highly stable resonance frequency with significantly weaker fluctuations (standard deviation $\approx 13.7$ MHz, a $34\times$ reduction from the open-loop scenario) even in the presence of additional perturbation ($\xi=7.55$). Figure 2(d) visualizes the time-domain data as histograms (bin size =5=5 MHz). Notably, even in the free-running case, the resonance frequency histogram does not exhibit a well-defined peak.